Speed for half journey is 30m/s and speed for remaining half journey is60 m/s then what is average speed for whole journey ?

Grade:12

1 Answers

Aman Kashyap
6 years ago
Let the total distance it covers be 2xTime taken to cover the first half of distance t1 = x/30Time taken to cover the next half of distance t2 = x/40Total time taken = t1 + t2 = x/30 + x/60 = x/20Average speed = total distance/total time taken = 2x/(x/20) = 40m/s